OpenStudy (anonymous):

4x^4 -15x^2 - 4 is divided by x -2

OpenStudy (anonymous):

4 0 -15 0 -4 2 8 16 2 4 --------------------- 4 8 1 2 0\[4x^3+8x^2+x+2\]If u like my answer ,click good answer,ty

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

quotient will be degree 3 Ax^3 + Bx^2 + Cx +D hence 4x^4 -15x^2 - 4 = (x -2)(Ax^3 + Bx^2 + Cx +D) [expand RHS] 4x^4 -15x^2 - 4 =Ax^4 + Bx^3+ Cx^2 +Dx -2Ax^3 -2Bx^2 -2Cx - 2D [group and factor] 4x^4 -15x^2 - 4 =Ax^4 +(B-A)x^3 +(C-2B)x^2 +(D-2C)x - 2D [equate coefficients] 4x^4 = Ax^4 A=4 (B-A) = 0 [no degree 3 coefficients] so B = 4 (C-2B) =-15 C= -7 - 2D = -4 D = 2 Hence answer = (4x^3 + 4x^2 -7x +2)

OpenStudy (anonymous):

(4x^3 + 4x^2 -7x +2) is wrong answer http://www.wolframalpha.com/input/?i=%284x%5E4+-15x%5E2+-+4+%29%C3%B7%28x+-2%29